### v1

@@ 1 Introduction
QPic has been formulated over two decades of research, pioneered by [1] and developed together with many others [32, 31, 34, 22, 24, 12, 17]. It has initially gained significance in academia, becoming an integral part of graduate and undergraduate courses at institutions worldwide. For over a decade, it has been taught across multiple courses in the departments of computer science, software engineering, and mathematics at the University of Oxford. It then has been incorporated into programs at other universities. [^1] Including the University of Amsterdam, University of Edinburgh, University of Cambridge, and Indiana University Bloomington. Additionally, it has been a central topic at international summer schools, conferences, and tutorials, demonstrating its growing influence across diverse academic and research settings globally. Graduate courses rooted in QPic have been taught at the University of Oxford for the past decade. Presently, five courses at the University of Oxford teach and examine QPic: the Quantum Processes and Computation graduate course in Computer Science, the Quantum Software graduate course (the successor to the former Categorical Quantum Mechanics course) in Computer Science, the Quantum Computing graduate course in Software Engineering, the Quantum Computing course in the Department for Continuing Education, and the Introduction to Quantum Information graduate course in Mathematics. Courses over the past year at other universities with multiple lectures of QPic in the curriculum include the Full-stack Quantum Computing Master’s in Quantum Computer Science course at the University of Amsterdam, the Introduction to Quantum Programming and Semantics undergraduate course at the University of Edinburgh, the Advanced Topics in Category Theory graduate course at the University of Cambridge, and the Programming Quantum Computers graduate course in summer 2023 at Indiana University Bloomington. Additionally, it has been the central topic of the Picturing Quantum Weirdness summer school at the University of Gdańsk in 2023 in Poland; a school at the Abdus Salam International School of Theoretical Physics (ICTP) in Trieste, Italy in 2023; two tutorials at the IEEE International Quantum Computing and Engineering Conference in 2023 in the US and in 2024 in Canada; a tutorial at the Quantum Initiative Rhineland-Palatinate (QUIP) International Winter School on Quantum Machine Learning 2024 at the Fraunhofer Institute for Industrial Mathematics ITWM in Germany; and a dozen lectures for the upcoming E-learning African International School on Quantum Science and Technology (ELAIS-QST) 2024/25 organized by the Kwame Nkrumah University of Science and Technology in Ghana. Note that this list does not include the growing implementation of QPic in the US or the increasing interest it is generating in other countries worldwide.
@@ 3.4 Whether QPic ensures the content is within the ability of students in this age range
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Upon reviewing the raw data, 20 participants achieved a distinction ($>70$), while 7 participants attained a merit (60-69), and 9 participants scored within the pass range ($50-59$), with 3 participants falling into the near fail category ($40-49$). Furthermore, 15 participants received a fail grade ($0-39$). However, two canary questions were included in the exam, a practice that has been implemented at the Oxford Computer Science exam tradition for over a decade. These questions were designed serve as indicators of participants’ overall understanding, with successful passing dependent on their successful completion. As detailed in Appendix D, canary questions acted as a critical benchmark in our assessment, typically setting a standard that ensures comprehension of the fundamental concepts or skills necessary for successful completion of the training. Therefore, those who were unable to solve these specific questions would not have been able to achieve passing grades, regardless of their performance on other sections of the exam. Figure 6 shows the final mark distribution, for all scripts and for scripts passing the canary test respectively. They are plotted in 10-mark buckets. Conditional on the canary test, the distribution shifts to higher marks, with no scripts marked below 10 and very few scripts marked below 20.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.6 Exam scores and demographic variables
[…]
@@ 6.5 Acknowledgements
We would like to extend our gratitude to the participating students, their supportive teachers, schools and families. Special thanks to Anna Pearson and Yoshi Yonezawa for their valuable comments through proofreading, and the tutors who assisted tutorials during the experiment. We are thankful to Sarah Meng Li for sharing her notes verifying quantum teleportation. We are also grateful to Karen Clayton for her help with logistical organisation.
@@ Appendix A A Crash Course on QPic
$\begin{pmatrix}0&0&0&1\\ 0&1&0&0\\ 0&0&1&0\\ 1&0&0&0\end{pmatrix}$
@@ Appendix E Marking criteria
The marking criteria outlined in Table E below are meticulously designed to target key assessment areas, aligning closely with the course’s learning objectives and expected outcomes. These criteria provide a clear and structured framework for evaluating student performance, ensuring that assessments accurately reflect the core competencies and skills intended for development in conjunction with discipline-specific criteria, and should be viewed as guidance on the overall standards expected at different grade bands, aligning with the taught postgraduate generic marking criteria used at the University of Oxford, a 0-100% grading structure in line with the current university regulations.
[…]
=1mm {tabu} —X[7.5, c]—X[7.5c]—X[7.5c]—X[10.5c]—   Distinction >>>70
@@
Advanced, in-depth, authoritative, full understanding of key ideas. Originality of the solutions, legitimacy of chain of reasoning in the answers provided. Complex work and key problems solved. Correct application of concepts and techniques (e.g. the appropriate use of diagrammatic rewriting rules), the ability to use proper terminology (e.g. “square-popping”, “leg-chopping”) Coherent and compelling work. Logical and concise presentation. The solution drawn/written in a clear and unambiguous way (e.g. the proper use of notations, the difference between “quantum” and “ordinary” diagrams). (90-100) insightful work displaying in-depth knowledge. Outstanding work, independent thought, highest standards of problem solving
(80-89) insightful work displaying in-depth knowledge. Good quality of work, independent thought (70-79) thoughtful work displaying in-depth knowledge, good standards of problem solving
@@
In-depth understanding of key ideas with evidence of some originality Key problems solved. Correct application of most concepts, techniques, and correct use of terminology Coherent work, logically presented. Clear solutions (65-69) thoughtful work displaying good knowledge and accuracy. Evidence for the ability to solve problems
(60-64) work displays good knowledge, some evidence for problem solving
@@
Understanding of some key ideas with evidence of ability to reflect critically Some key problems solved. Correct application of some concepts, techniques, and terminology Competent work in places but lacks coherence (55-59) work displays some understanding in most areas, but standard of work is variable
(50-54) work displays knowledge and understanding of some areas, but some key problems are not solved
@@
Lack of understanding Key problems misunderstood/unanswered, limited/incorrect application of concepts, techniques and terminology Work is confused and incoherent (33-39) incomplete answers with some superficial knowledge
(20-32) some attempt to write something relevant but many flaws (0-19) serious errors, irrelevant answers

### v2

@@ 1 Introduction
QPic has been formulated over two decades of research, pioneered by [1] and developed together with many others [32, 31, 34, 22, 24, 12, 17]. It has initially gained significance in academia, becoming an integral part of graduate and undergraduate courses at institutions worldwide. For over a decade, it has been taught across multiple courses in the departments of computer science, software engineering, and mathematics at the University of Oxford. It then has been incorporated into programs at other universities [^1] Including the University of Amsterdam, University of Edinburgh, University of Cambridge, and Indiana University Bloomington. Additionally, it has been a central topic at international summer schools, conferences, and tutorials, demonstrating its growing influence across diverse academic and research settings globally. Graduate courses rooted in QPic have been taught at the University of Oxford for the past decade. Presently, five courses at the University of Oxford teach and examine QPic: the Quantum Processes and Computation graduate course in Computer Science, the Quantum Software graduate course (the successor to the former Categorical Quantum Mechanics course) in Computer Science, the Quantum Computing graduate course in Software Engineering, the Quantum Computing course in the Department for Continuing Education, and the Introduction to Quantum Information graduate course in Mathematics. Courses over the past year at other universities with multiple lectures of QPic in the curriculum include the Full-stack Quantum Computing Master’s in Quantum Computer Science course at the University of Amsterdam, the Introduction to Quantum Programming and Semantics undergraduate course at the University of Edinburgh, the Advanced Topics in Category Theory graduate course at the University of Cambridge, and the Programming Quantum Computers graduate course in summer 2023 at Indiana University Bloomington. Additionally, it has been the central topic of the Picturing Quantum Weirdness summer school at the University of Gdańsk in 2023 in Poland; a school at the Abdus Salam International School of Theoretical Physics (ICTP) in Trieste, Italy in 2023; two tutorials at the IEEE International Quantum Computing and Engineering Conference in 2023 in the US and in 2024 in Canada; a tutorial at the Quantum Initiative Rhineland-Palatinate (QUIP) International Winter School on Quantum Machine Learning 2024 at the Fraunhofer Institute for Industrial Mathematics ITWM in Germany; and a dozen lectures for the upcoming E-learning African International School on Quantum Science and Technology (ELAIS-QST) 2024/25 organized by the Kwame Nkrumah University of Science and Technology in Ghana. Note that this list does not include the growing implementation of QPic in the US or the increasing interest it is generating in other countries worldwide., as well as a wide range of outreach initiatives involving high school students and teachers [^2] Including the KYMA program, aimed at introducing Greek high school students to the world of Quantum Information, through a collaboration between secondary and university instructors, supported by the Hellenic Governmental Authorities and Hellenic Mathematical Society; the Quantum Physics in Pictures course taught by Muhammad Hamza Waseem and hosted by Khwarizmi Science Society, Pakistan; a two-day summer camp for high schoolers hosted at the University of Dalhousie; the 2025 Intro to Quantum Teacher Workshop for middle and high school teachers ahead of QSEEC; the QIST in the CSU Faculty Workshop..
@@ 3.4 Whether QPic ensures the content is within the ability of students in this age range
We also considered the feeling of preparedness for exam questions. This feeling negatively correlated only with physics grade, suggesting that participants who felt less prepared tended to have higher physics grades, while those who felt more prepared tended to have lower physics grades. The question regarding feeling of preparedness was correlated negatively with the final exam difficulty, as might be expected, [r(59)=0.77,(p<.001)r(59)=-0.77,(p<.001)]. This disparity between perceived preparedness and actual exam performance in A levels needs further investigation to understand the underlying factors at play, but it is possibly just the result of students deciding to devote more or less effort to the quantum course vs. their A levels.
@@
There were positive correlations between gender, Physics, and Mathematics grades. The correlations indicate that as we go from females (labeled with a 1) to males (labeled with 2), performance in Physics and Mathematics improved. Independent samples t-test confirmed significant results for both Mathematics grades [t(36.033)=2.577t(36.033)=-2.577, $p<.001$, CI(-.911, -.128)] and Physics grades [t(31.966)=2.664,df(p<.001)t(31.966)=-2.664,df(p<.001), CI(-.911, -.128)], with male participants scoring higher than female counterparts in both subjects. There is clearly work that needs to be done, to understand the role of gender in determining the performance levels in these subjects. Note, this trend is consistent with the literature on gender differences in STEM fields, with the general finding being that males tend to outperform females in physics and mathematics [58, 48, 114]. Such an outcome might be influenced by various socio-cultural factors and educational practices, as discussed by [58]. Interestingly, this trend does not extend to English grades, indicating that gender may have a differential impact on performance across different subjects.
Another noteworthy finding was the strong difference in all of Mathematics, Physics, and English A level grades, between participants with families in owned (higher scores) vs. rented accommodation. As we go from renters to homeowners, participants from homeowner families tended to achieve higher grades in Mathematics, Physics, and English as highlighted by the negative correlations. Possible explanations include stronger academic motivation among students from more advantaged family backgrounds, along with potential disparities in access to resources or support systems. For example, students from homeowner families may receive additional assistance or interventions to improve their skills in these subjects. Using independent samples t-tests to compare scores from the three subjects, we observed for Mathematics t(20.303)=4.652,p<.001t(20.303)=-4.652,p<.001, CI(.499 - 1.258), Physics t(24.979)=2.575,p<.001t(24.979)=-2.575,p<.001, CI(.120 - .974)], and English t(50)=2.577,p.01t(50)=2.577,p-.01, CI(.203, 1.635). Note, the family status variable indicated differences in additional commitments too: we created a variable by adding 0/1 indicator variables, across four categories: additional commitments regarding an exam, a part time job, an internship, some support activity, or some leisure activity (e.g., a club commitment). However, there was no difference in overall additional commitments between students from home owning families vs. families renting their accommodation ($p=.87$).
@@
Mathematics grades negatively correlated with those involved in additional camp or leisure activities ($r=.269,p<.045$), but not with any other variable, though it is unclear why leisure and camp commitments may impact Mathematics grades specifically, but not Physics and English grades.
@@
We asked participants to evaluate how engaging they found the tutors’ communication and teaching style. The majority rated it as either good (40.7%) or very good (29.6%). Additionally, there was a highly significant correlation between this evaluation and the item “the course motivated me to pursue a STEM career.” This indicates that the experiment and teaching style significantly increased participants’ motivation to pursue a STEM career, demonstrating a sizeable positive impact ($r(54)=.419,p=.001$) on participants – this result attests to the achievement of one of the primary goals of the experiment in supporting young people entering the STEM fields. Significant correlations between this item and ‘the course strengthened my motivation’ ($r(54)=.261,p=.02$) further validate this outcome.
Interestingly, the number of hours spent studying per week showed no correlation with most variables. One exception was the question about how much one enjoyed the course ($r(57)=.307,p=.02$).
@@
The key dependent variable of the present study concerns the overall, final score in the quantum exam, following the training. With a pass mark of 50%, minimally we are interested in the extent to which there is evidence of above chance performance. This proved to be the case, as shown by a one-sample t-test against chance, $t(53)=2.166,p=.035$, with a mean exam score of $57.93$ and standard deviation of $26.895$. The distributional characteristics of the exam scores in Figure 6 shows a broad range of individuals, in terms of performance, from a sizable group of pupils with solid, over 70%, to many who performed well below chance.
@@
Upon reviewing the raw data, 20 participants achieved a distinction (>70>70), while 7 participants attained a merit (60-69), and 9 participants scored within the pass range (505950-59), with 3 participants falling into the near fail category (404940-49). Furthermore, 15 participants received a fail grade (0390-39). However, two canary questions were included in the exam, a practice that has been implemented at the Oxford Computer Science exam tradition for over a decade. These questions were designed serve as indicators of participants’ overall understanding, with successful passing dependent on their successful completion. As detailed in Appendix D, canary questions acted as a critical benchmark in our assessment, typically setting a standard that ensures comprehension of the fundamental concepts or skills necessary for successful completion of the training. Therefore, those who were unable to solve these specific questions would not have been able to achieve passing grades, regardless of their performance on other sections of the exam. Figure 6 shows the final mark distribution, for all scripts and for scripts passing the canary test respectively. They are plotted in 10-mark buckets. Conditional on the canary test, the distribution shifts to higher marks, with no scripts marked below 10 and very few scripts marked below 20.
@@
We statistically confirmed the impression that Coecke assigned higher marks than Kissinger (overall one-way repeated measures ANOVA, $F(2,18)=19.3,p<.001$ and corresponding pairwise comparison between Coecke and Kissinger $t(53)=4.57,p<.001$; note the degrees of freedom are different between the omnibus ANOVA and the paired-samples comparison, because of missing values concerning Gogioso’s marking).
@@
  • The developmental dynamics within the studied age range (16 to 18 years), it is important to note that all participants were recruited within this limited range, with an average mean of 16.84 years and a standard deviation of $0.58$. A correlation between age and exam scores did not reveal evidence for a reliable trend (r(54)=0.19r(54)=-0.19). An alternative approach to exploring the dependence of exam scores on age would be to treat the four different ages amongst our participants as a nominal variable and consider possible differences via a one-way, between participants ANOVA. As can be seen in Figure 11 below, there is a trend for lower exam scores, with participants who were 18 years old versus the younger ones (16, 17), but without any evidence for statistical reliability, $F(1,48)=1.45,p=.24$. Note, the differences in error bars might indicate a violation of the homogeneity of variance assumption, but this was not the case (using Levene’s test, $p=.78$). While we are reluctant to interpret a non-significant trend, it can be perhaps speculated that older participants had more varied and pressing responsibilities, such as A Level exams.
@@
  • Regarding location of residence, most participants (13/54) were from London, but there was a sizable number of participants with missing information (16/54). As 24 different locations were represented, we decided to aggregate into the following categories, for analyses: London, larger cities and towns with a university tradition (Cambridge, Bristol, Manchester etc.), and other cities and towns. Figure 11 below provides a visual impression of how scores differ, across the three types of location of residence. The trend for lower scores for participants from ‘other cities and towns’ was not significant, $F(2,35)=0.96,p=.39$ (note, the residual degrees of freedom are lower in this comparison, because of missing values). It is perhaps unsurprising, as for many locations we had only one participant, which precludes systematic analyses.
@@
Figure 14 shows means and error bars for average performance across the three categories. As can be seen, the means are nearly identical and a one-way between participants ANOVA failed to reach significance, $F(2,48)=0.32,p=.72$ (note, Levene’s test offered no evidence for violations of the assumption of homogeneity of variance, $p=.56$). Further examination shows that, while mean performance for state boarding and private independent participants was nearly identical to that from state and grammar school participants, we have some extremely high variability regarding private boarding participants (this is partly due to the low N in this category).
@@ 3.6 Exam scores and demographic variables
As can be seen, the only statistically reliable result concerned an increased motivation to pursue a STEM career, based on the enjoyment of the course ($r(39)=.78,p<.001$). Interestingly, there was no corresponding association between STEM career motivation and the actual exam result, with a correlation of nearly zero (r(39)=.002,p=.99r(39)=-.002,p=.99). Regarding gender, school type, and location (all coded as above), we considered the same set of correlations separately for the different levels of each variable, with one variable level examined at a time (this approach is an informal way of examining interactions between the original variables in the correlations analysis and gender, school type, and location; it can help obtain a first impression of the direction of any possible interactions, if they exist). Regarding gender, the correlation between enjoyment and STEM career motivation was slightly stronger for male participants ($r(19)=.85,p<.001$) than female participants ($r(19)=.70,p<.001$); no other associations emerged as notable in these subgroups (here and elsewhere, we are interested in associations concerning exam scores, motivation for STEM careers and the rest of the variables). Regarding school type, for state, grammar, and private and boarding schools, we observed correlations of .76 (N=21), .78 (N=10), and .98 (N=8), respectively (all significant). Additionally, a very high correlation emerged between motivation for STEM career and max study hours, for private and boarding schools ($r(5)=.92,p=.028$). Finally, an equivalent analysis for location offered broadly similar conclusions. For participants for whom we had no location information, were based in London, were based in larger cities and university towns, and were based in other cities and towns, the correlations between STEM career motivations and course enjoyment were, respectively, r(12)=.81(p=.002),r(10)=.65,p=.044,r(7)=26,p=.58r(12)=.81(p=.002),r(10)=.65,p=.044,r(7)=26,p=.58, and $r(10)=.92,p<.001$. No other correlations were significant.
@@ 6.5 Acknowledgements
We would like to extend our gratitude to the participating students, their supportive teachers, schools and families. Special thanks to Anna Pearson and Yoshi Yonezawa for their valuable comments through reviewing, and the tutors who assisted tutorials during the experiment. We are thankful to Sarah Meng Li for sharing her notes verifying quantum teleportation. We are also grateful to Karen Clayton for her help with logistical organisation.
@@ Appendix A A Crash Course on QPic
$\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix}$
@@ Appendix E Marking criteria
The marking criteria outlined in Table 15 below are meticulously designed to target key assessment areas, aligning closely with the course’s learning objectives and expected outcomes. These criteria provide a clear and structured framework for evaluating student performance, ensuring that assessments accurately reflect the core competencies and skills intended for development in conjunction with discipline-specific criteria, and should be viewed as guidance on the overall standards expected at different grade bands, aligning with the taught postgraduate generic marking criteria used at the University of Oxford, a 0-100% grading structure in line with the current university regulations.
Distinction >>70
@@
Advanced, in-depth, authoritative, full understanding of key ideas. Originality of the solutions, legitimacy of chain of reasoning in the answers provided. Complex work and key problems solved. Correct application of concepts and techniques (e.g. the appropriate use of diagrammatic rewriting rules), the ability to use proper terminology (e.g. “square-popping”, “leg-chopping”) Coherent and compelling work. Logical and concise presentation. The solution drawn/written in a clear and unambiguous way (e.g. the proper use of notations, the difference between “quantum” and “ordinary” diagrams). (90-100) insightful work displaying in-depth knowledge. Outstanding work, independent thought, highest standards of problem solving (80-89) insightful work displaying in-depth knowledge. Good quality of work, independent thought (70-79) thoughtful work displaying in-depth knowledge, good standards of problem solving
@@
In-depth understanding of key ideas with evidence of some originality Key problems solved. Correct application of most concepts, techniques, and correct use of terminology Coherent work, logically presented. Clear solutions (65-69) thoughtful work displaying good knowledge and accuracy. Evidence for the ability to solve problems (60-64) work displays good knowledge, some evidence for problem solving
@@
Understanding of some key ideas with evidence of ability to reflect critically Some key problems solved. Correct application of some concepts, techniques, and terminology Competent work in places but lacks coherence (55-59) work displays some understanding in most areas, but standard of work is variable (50-54) work displays knowledge and understanding of some areas, but some key problems are not solved
@@
Lack of understanding Key problems misunderstood/unanswered, limited/incorrect application of concepts, techniques and terminology Work is confused and incoherent (33-39) incomplete answers with some superficial knowledge (20-32) some attempt to write something relevant but many flaws (0-19) serious errors, irrelevant answers